OpenStudy (anonymous):

A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground? a) 2.88pi inches b) 5.78pi inches c) 11.56pi inches

OpenStudy (anonymous):

I tried to divide 80 by 26 and then multiplying that answer by 180, but it didn't work out so well...

OpenStudy (anonymous):

now we have to convert 80 degree to radians

OpenStudy (anonymous):

Divide it by pi

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no this time we are converting the other way \[80\times \frac{\pi}{180}\]

OpenStudy (anonymous):

1.395?

OpenStudy (anonymous):

that comes out to be \[\frac{4\pi}{9}\] radians

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

0.4pi

OpenStudy (anonymous):

yeah about that we are still not done though

OpenStudy (anonymous):

Do we multiply that by 180?

OpenStudy (anonymous):

no now we think, maybe draw a picture

OpenStudy (anonymous):



OpenStudy (anonymous):

we want \(a\) the arc length the radian measure is arc length over the radius so \[\frac{a}{26}=.4\pi=1.4\] so \[a=26\times 1.4\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=26*1.4 you got a choice that is close to 36.4?

OpenStudy (anonymous):

oh wait, i see all your answers have \(\pi\) in them

OpenStudy (anonymous):

\[a=26\times \frac{9}{4}\pi=11.5\pi\]

OpenStudy (anonymous):

Wow that makes alot more sense than it did before

OpenStudy (anonymous):

yeah sorry i messed up earlier

OpenStudy (anonymous):

Umm. There's another question. I have already answered it, but I want to make sure I'm doing this the right way

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

If the length of an arc is 12pi inches and the radius of the circle is 10 inches, what is the measure of the arc? a) 216 degrees b) 270 degrees c) 288 degrees

OpenStudy (anonymous):

I chose A

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

the radian measure is \[\frac{12\pi}{10}=1.2\pi\]

OpenStudy (anonymous):

then \[1.2\pi\times \frac{180}{\pi}=1.2\times 180=216\]

OpenStudy (anonymous):

yeah you got it !

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw now i am listening to your your tube

OpenStudy (anonymous):

I feel pretty good now. Finally! Lol!

OpenStudy (anonymous):

Take your time.

OpenStudy (anonymous):

what should i be looking for ?>

OpenStudy (anonymous):

Me dropping my music and then picking it up on the left side of the screen somewhere in the middle of the song

OpenStudy (anonymous):

Lol!

OpenStudy (anonymous):

I should be the last person on the row of trumpets. On the outside since I'm leader

OpenStudy (anonymous):

i will look again right or left side?

OpenStudy (anonymous):

Left

OpenStudy (anonymous):

You should see white paper

OpenStudy (anonymous):

on the floor and me picking it up

OpenStudy (anonymous):

ok looking now

OpenStudy (anonymous):

You see it. I mean me Lol!

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

you are hidden mostly but i did see you pick up the sheet!

OpenStudy (anonymous):

Yeah It's a little embarrassing

OpenStudy (anonymous):

you got more math questions are are you ok? i gotta run soon

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

I had to grab something to eat

OpenStudy (anonymous):

eating is good you got any more before i run?

OpenStudy (anonymous):

Yeah